OpenStudy (anonymous):

A student needs to make a circular cardboard piece with an area between 154 square inches and 616 square inches. The function f(r) relates the area of the cardboard piece, in square inches, to the radius r in inches. Which of the following shows a reasonable domain for f(r)? 7 < r < 14 154 < r < 616 All positive integers less than 7 All positive integers between 14 and 154 @Godlovesme

OpenStudy (godlovesme):

the area depends on the radius right

OpenStudy (anonymous):

yes

OpenStudy (godlovesme):

so the dependent will be the area --> f(r) --> range independent will be radius --> r --> domain

OpenStudy (godlovesme):

confused?

OpenStudy (anonymous):

yes, kind of.

OpenStudy (anonymous):

I was thinking the answer would be A, but I'm not positive

OpenStudy (jhannybean):

First define the area of a circle:\[A=\pi r^2\]

OpenStudy (jhannybean):

How you're given \(A_1 = 616~,~ A_2 = 154\)

OpenStudy (jhannybean):

Now*.

OpenStudy (dtan5457):

To get your radius. \[\sqrt{\frac{ A }{ \pi }} \]

OpenStudy (jhannybean):

So to find the radius of each, as @dtan5457 mentioned, \[r^2 = \frac{A}{\pi} \implies r = \sqrt{\frac{A}{\pi}}\]

OpenStudy (jhannybean):

Now find \(r_1~,~r_2\) with respect to \(A_1~,~A_2\)

OpenStudy (dtan5457):

\[A=\pi r^2\] This is the regular formula for area of a circle, so that's why you would use the first formula as a vice versa to get back to your original radius.

OpenStudy (jhannybean):

\[r_1 = \sqrt{\frac{A_1}{\pi}} = \sqrt{\frac{616}{\pi}}\]\[r_2 = \sqrt{\frac{A_2}{\pi}} = \sqrt{\frac{154}{\pi}}\]

OpenStudy (dtan5457):

Your two areas were originally 154 square inches and 616 square inches. Put a square root on your calculator, and divide those numbers by the pi key.

OpenStudy (dtan5457):

@Jhannybean already did the hard part for you